# ZeroQuant(4+2): Redefining LLMs Quantization with a New FP6-Centric Strategy for Diverse Generative Tasks

### 14 Dec 2023 v1

@@ 4 Sweet Spot Solution: FP6
where is the original full precision weight, symbolizes the quantization function, is the scaling factor, and is the FP6 number. The scaling factor is computed by: thereby ensuring optimal use of ’s range without compromising on precision. Please see Section 5 for additional customizations in FP6.
@@ 4.1 Results of FP6 and FP5 on all tasks
FP6 on CGQ Performance Comparable to Baseline. As indicated in Tables 6 and 6, FP6 employing CGQ (highlighted in dark orange) successfully attains accuracy levels equivalent to those of full precision. Additionally, FP5 results were evaluated using CGQ, revealing a disparity compared to the baseline.
@@
FP6 Performance and Robustness. FP6 quantization, especially with CGQ, demonstrates a significant advancement, nearly matching the FP16 baseline across various tasks and models. This quantization method not only narrows the performance gap seen in FP5 and INT4 but also shows robustness in handling different tasks. The robustness is further accentuated when comparing CGQ and FGQ within the FP6 framework (as there is little difference between CGQ and FGQ), where FP6 with CGQ consistently maintains high performance close to baseline, indicating its effectiveness and stability across different scenarios. Moreover, FP6 is robust to quantization algorithms: either RTN or GPTQ results in similar results, particularly for CodeLLaMA-34B, as shwon in Table 6.
@@ 5.3 System evaluation
We conducted an evaluation of the latest GPU kernels across various weight-only quantization techniques, focusing on their system performance. This kernel-level assessment was carried out on the NVIDIA A100-40GB platform, running Linux 5.3.18 and CUDA 11.8. Our primary focus was on the performance of feed-forward (FFN) layers within the LLaMA models, specifically during the token generation phase, as detailed in [57]. Comprehensive data on matrix shapes and kernel latency is available in Appendix B. We employed cuBLAS [48] as our benchmark for non-quantized performance (W16A16). We also included cutting-edge kernel support for F INT4 FGQ quantization (W4A16) from TensorRT-LLM [49] for comparative analysis.121212The support block-size 64 and 128, we chose 128.
@@ 6 Discussion and Conclusion
Xiaoxia Wu led model quality measurement and paper writing. Haojun Xia, contributing to system section writing, designed the FP6 acceleration strategy and system support. Stephen Youn, Zhen Zheng, and Shiyang Chen handled DeepSpeed integration. Arash Bakhtiari and Michael Wyatt managed framework testing and external integration. Yuxiong He and Olatunji Ruwase provided key discussions. Leon Song is the overall leader for algorithm and system design. Zhewei Yao initiated the project and led the FP6/5 format design.
@@ Appendix A Background of Quantization
Throughout this work, we focus on post-training quantization (PTQ), i.e., no or minimal training effort is applied after quantization, for which large accuracy degradation usually exhibits for coarse-grained quantization (per matrix/tensor) due to their large quantization error. As such, we focus on fine-grained quantization. Particularly, we use the per-row quantization (one row of the weight matrix or one token for the activation) from [68] as our coarsest-grained quantization method, and we use block-k quantization (for every k elements, they have their own scaling factor and/or zero point) as our finer-grained quantization scheme.

### 18 Dec 2023 v2

@@ 4 Sweet Spot Solution: FP6
where is the original full precision weight, symbolizes the quantization function, is the scaling factor, and is the FP6 number. The scaling factor is computed by: thereby ensuring optimal use of ’s range without compromising on precision. Please see Section 5 for additional customizations in FP6. Similar settting is defined for F5.
@@ 4.1 Results of FP6 and FP5 on all tasks
FP6 Robustness. FP6 quantization, especially with CGQ, demonstrates a significant advancement, nearly matching the FP16 baseline across various tasks and models. This quantization method not only narrows the performance gap seen in FP5 and INT4 but also shows robustness in handling different tasks. The robustness is further accentuated when comparing CGQ and FGQ within the FP6 framework (as there is little difference between CGQ and FGQ), where FP6 with CGQ consistently maintains high performance close to baseline, indicating its effectiveness and stability across different scenarios. Moreover, FP6 is robust to quantization algorithms: either RTN or GPTQ results in similar results, particularly for CodeLLaMA-34B, as shwon in Table 6.
@@ 5.3 System evaluation
We conducted an evaluation of the latest GPU kernels across various weight-only quantization techniques, focusing on their system performance. This kernel-level assessment was carried out on the NVIDIA A100-40GB platform, running Linux 5.3.18 and CUDA 11.8. Our primary focus was on the performance of feed-forward (FFN) layers within the LLaMA models, specifically during the token generation phase, as detailed in [57]. Comprehensive data on matrix shapes and kernel latency is available in Appendix B. We employed cuBLAS [48] as our benchmark for non-quantized performance (W16A16). We also included cutting-edge kernel support for F INT4 FGQ quantization (W4A16) from TensorRT-LLM [49] for comparative analysis.121212Between the supported block-size: 64 and 128, we chose 128.
@@ 6 Discussion and Conclusion
Xiaoxia Wu led model quality measurement and paper writing. Haojun Xia, contributing to system section writing, designed the FP6 acceleration strategy and system support. Stephen Youn, Zhen Zheng, and Shiyang Chen handled DeepSpeed integration. Arash Bakhtiari and Michael Wyatt managed framework testing and external integration. Reza Yazdani Aminabadi, Yuxiong He and Olatunji Ruwase provided key discussions. Leon Song is the overall leader for algorithm and system design. Zhewei Yao initiated the project and led the FP6/5 format design.
@@ Appendix A Background of Quantization
Throughout this work, we focus on post-training quantization (PTQ), i.e., no or minimal training effort is applied after quantization, for which large accuracy degradation usually exhibits for coarse-grained quantization (per matrix/tensor) due to their large quantization error. Particularly, we use the per-row quantization (one row of the weight matrix) from [68] as our coarsest-grained quantization method, and we use block-k quantization (for every k elements, they have their own scaling factor and/or zero point) as our finer-grained quantization scheme.